# Thermodynamic Consistent Description of Compact Stars of Two Interacting Fluids: The Case of Neutron Stars with Higgs Portal Dark Matter

### v1

@@ 1 Introduction
Among others, DM could influence the physical properties and dynamics of compact stars, such as neutron stars, thus leading to a deeper understanding of these objects and the broader dark sector [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Indirect detection possibilities, such as neutrino and gamma-ray signals from inelastic DM annihilations near neutron stars, further increase the potential for uncovering DM properties [37]. Neutron stars, acting as natural DM detectors, provide crucial insights into DM properties through their sensitivity to DM interactions and annihilation processes within their cores [25, 38]. Furthermore, the concept of supramassive dark objects broadens insights into DM’s influence on stellar evolution, and white dwarfs mixed with DM may collapse into neutron stars, affecting the overall stellar population [35, 39, 40, 41].
@@
Gravitational wave astronomy is another possible tool to detect DM within neutron stars. Observations from experiments such as LIGO (Laser Interferometer Gravitational-Wave Observatory) and Virgo, including events like GW170817, provide insights into the structure of neutron stars [42, 43]. Non-radial oscillation modes in DM-admixed neutron stars suggest that these oscillations could serve as indicators of DM presence [44, 30, 45, 28, 46]. Additionally, the study of tidal Love numbers—parameters that quantify a star’s tidal response to external gravitational fields—enhances our understanding of the composition of neutron stars [47, 48, 49, 50] and allows for elucidating the possible existence and nature of DM.
@@
Specific DM candidates, including axions, dark photons, and ultralight particles, exhibit unique astrophysical signatures that can be probed through neutron star studies. Axions, for example, may address the cusp-core problem in galactic halos and significantly influence neutron star cooling processes [51, 52, 53]. Dark photons impact neutron star stability and could generate observable signals detectable through various astrophysical observations [54, 55, 56, 39]. Ultralight dark photons and millicharged particles extend DM models with significant implications for neutron star stability and dynamics [54, 55, 57], while studies suggesting neutron decay into a dark Dirac fermion and a dark photon could explain discrepancies in neutron lifetime measurements, offering new pathways for DM detection [56, 58]. Exotic compact objects, such as mirror neutron stars and boson stars, offer valuable insights into DM’s astrophysical role. Mirror matter within neutron stars could produce distinctive gravitational wave signatures, differentiating these stars from their ordinary counterparts [59, 60, 4]. Self-interacting DM models, particularly those involving the Higgs portal and muonic forces, are investigated using astrophysical data to verify their viability [61, 62, 63]. Boson stars, formed from bosonic DM with repulsive self-interactions, present alternative compact objects that challenge traditional neutron star models [64, 65, 66, 67, 68, 69]. Studies on fermion-boson stars aim to reconcile observations of compact stars with complex DM interactions, bridging theoretical predictions with astronomical data [10, 65, 70, 71, 72]. Moreover, asymmetric bosonic DM, mediated by vector fields coupled to scalar particles, may alter neutron stars’ mass, radius, and tidal deformability [13, 73, 74, 16, 75, 76]. Non-annihilating particle DM can slowly accumulate within neutron stars, potentially leading to the formation of low-mass black holes, a scenario supported by LIGO-Virgo observations [77, 41, 40, 78, 79, 3]. DM may also influence neutron star heating and cooling, with density-dependent EoS models showing that DM admixtures impact thermal evolution and axion emissions from protoneutron stars suggest possible indirect detection of these particles [74, 52, 53, 80, 81, 82, 83, 84, 85, 86, 87]. These cooling mechanisms, governed by energy loss processes such as neutrino emission, can constrain axion properties and their coupling strengths to ordinary matter, offering a window into physics beyond the Standard Model [80, 52, 53].
@@ 2 Structure Equations: Two Approaches
$\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,,$ (2.4)
@@
$\frac{dp(r)}{dr}$ $=$ $-\left(\epsilon(r)+p(r)\right)\frac{d\nu(r)}{dr}\,,$ (2.5)
@@
$\frac{dM(r)}{dr}$ $=$ $4\pi r^{2}\epsilon(r)\,,$ (2.6)
@@
$\mu=\frac{\epsilon+p}{\rho}\,,$ (2.7)
or, alternatively, it can be derived from the energy density as the derivative with respect to the number density 111 For nucleonic matter, one can solve the structure equations based on the baryon chemical potential $\mu_{\rm N}$ by considering the baryon/nucleon number density $\rho_{\rm N}$, which includes both neutrons and protons $\mu_{\rm N}=\frac{\epsilon+p}{\rho_{\rm N}}\,,\quad\mu_{\rm N}=\mu_{n}+\mu_{p}\,,\quad\rho_{\rm N}=\rho_{n}+\rho_{p}\,.$ (2.8) This approach allows for a unified treatment of the matter inside neutron stars, accounting for the contributions from different particle species.
$\mu=\frac{d\epsilon}{d\rho}\,.$ (2.9)
@@
$\nabla_{\mu}T^{\mu\nu}_{\rm N}=-Q^{\nu},\quad\nabla_{\mu}T^{\mu\nu}_{\rm DM}=Q^{\nu}\,,$ (2.12)
@@
$\frac{d\mu_{i}(r)}{dr}=-\mu_{i}(r)\frac{d\nu(r)}{dr}\,,$ (2.13)
$\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,,$ (2.14)
$\frac{dM(r)}{dr}=4\pi r^{2}\epsilon(r)\,,$ (2.15)
@@
$\frac{dp_{i}(r)}{dr}=-(\epsilon_{i}(r)+p_{i}(r))\frac{d\nu(r)}{dr}\,,$ (2.16)
$\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,.$ (2.17)
$\frac{dM(r)}{dr}=4\pi r^{2}\epsilon(r)\,,$ (2.18)
@@ 2.1 Dimensionless structure equations for the interacting DM-nucleonic system
$M^{\prime}=\frac{M}{M_{\rm L}}\,,\quad M_{\rm L}=\frac{M_{\rm Pl}^{3}}{m_{\rm f}^{2}}\,,\quad r^{\prime}=\frac{r}{R_{\rm L}}\,,\quad R_{\rm L}=\frac{M_{\rm Pl}}{m_{\rm f}^{2}}\,,\quad\mu^{\prime}=\frac{\mu}{m_{f}}\,,$ (2.19)
@@
$\frac{d\nu(r^{\prime})}{dr^{\prime}}$ $=$ $\frac{\left(M^{\prime}(r^{\prime})+4\pi r^{\prime 3}p^{\prime}(r^{\prime})\right)}{r^{\prime}\left(r^{\prime}-2M^{\prime}(r^{\prime})\right)}\,,$ (2.20)
$\frac{dM^{\prime}(r^{\prime})}{dr^{\prime}}$ $=$ $4\pi r^{\prime 2}\epsilon^{\prime}(r^{\prime})\,,$ (2.21)
$\frac{d\mu^{\prime}_{\rm N}(r^{\prime})}{dr^{\prime}}$ $=$ $-\mu^{\prime}_{\rm N}(r^{\prime})\frac{d\nu(r^{\prime})}{dr^{\prime}}\,,$ (2.22)
$\frac{d\mu^{\prime}_{\rm DM}(r^{\prime})}{dr^{\prime}}$ $=$ $-\mu^{\prime}_{\rm DM}(r^{\prime})\frac{d\nu(r^{\prime})}{dr^{\prime}}\,.$ (2.23)
@@ 3 The Higgs Portal Fermionic Dark Matter Model
$\mathcal{L}=$ $\ \bar{\psi}\left[\gamma^{\mu}\left(i\partial_{\mu}-g_{v}V_{\mu}-g_{\rho}\tau\cdot b_{\mu}\right)-\left(M_{\rm N}+g_{s}\phi\right)\right]\psi$
$+\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m_{s}^{2}\phi^{2}-\frac{1}{3}g_{2}\phi^{3}-\frac{1}{4}g_{3}\phi^{4}$
$-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-\frac{1}{4}b_{\mu\nu}b^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}b_{\mu}b^{\mu}\,,$ (3.1)
@@
$V_{\mu\nu}$ $=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}\,,$ (3.2)
$b_{\mu\nu}$ $=\partial_{\mu}b_{\nu}-\partial_{\nu}b_{\mu}\,.$ (3.3)
@@
$\mathcal{L}_{\text{\rm DM}}=$ $\ \bar{\chi}\left(i\gamma^{\mu}\partial_{\mu}-M_{\chi}+yh\right)\chi+\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}M_{h}^{2}h^{2}+f\frac{M_{\rm N}}{v}\bar{\psi}h\psi\,,$ (3.4)
where $h$ is the Higgs field, $y$ is the Yukawa coupling between DM and the Higgs, $f$ represents the nucleon-Higgs form factor  [^2] We fix the parameters to $M_{h}=125$ GeV [112], $y=0.07$ and $f=0.35$ as a reasonable choice, see e.g. Ref. [7]., $M_{\rm N}$ and $M_{\chi}$ are the nucleon and $\chi$ masses, respectively, and $v$ is the vacuum expectation value of the Higgs field. The interaction term $f\frac{M_{\rm N}}{v}\bar{\psi}h\psi$ accounts for the coupling between nucleons and the Higgs field [113].
@@
$M_{\rm N}^{*}$ $=M_{\rm N}+g_{s}\phi_{0}-\frac{fM_{\rm N}}{v}h_{0}\,.$ (3.6)
As for the mesons, the field equations of motion follow from the Euler–Lagrange equations, where in the mean-field approximation, the meson fields are replaced by their classical expectation values (denoted by the subscript 00)[7, 107]
$\phi_{0}$ $=\frac{1}{m_{s}^{2}}\left(-g_{s}\langle\bar{\psi}\psi\rangle-g_{2}\phi_{0}^{2}-g_{3}\phi_{0}^{3}\right)\,,$ (3.7)
$V_{0}$ $=\frac{g_{v}}{m_{V}^{2}}\langle\psi^{\dagger}\psi\rangle=\frac{g_{v}}{m_{V}^{2}}(\rho_{p}+\rho_{n})\,,$
$b_{0}$ $=\frac{g_{\rho}}{m_{\rho}^{2}}\langle\psi^{\dagger}\tau_{3}\psi\rangle=\frac{g_{\rho}}{m_{\rho}^{2}}(\rho_{p}-\rho_{n})\,,$
$h_{0}$ $=\frac{y\langle\bar{\chi}\chi\rangle+f\frac{M_{\rm N}}{v}\langle\bar{\psi}\psi\rangle}{M_{h}^{2}}\,.$
@@
$\rho$ $=\langle\psi^{\dagger}\psi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}d^{3}k\,,$ (3.8)
$\rho_{\rm N}^{s}$ $=\langle\bar{\psi}\psi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}\frac{M_{\rm N}^{*}}{\sqrt{M_{\rm N}^{*2}+k^{2}}}d^{3}k\,,$ (3.9)
$\rho_{\text{\rm DM}}^{s}$ $=\langle\bar{\chi}\chi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}\frac{M_{\chi}^{*}}{\sqrt{M_{\chi}^{*2}+k^{2}}}d^{3}k\,,$ (3.10)
@@
$M_{\chi}^{*}$ $=M_{\chi}-yh_{0}\,,$ (3.11)
[…]
$\mu_{p}$ $=g_{v}V_{0}+g_{\rho}b_{0}+\sqrt{k_{p}^{2}+(M_{\rm N}^{*})^{2}}\,,$ (3.14)
$\mu_{n}$ $=g_{v}V_{0}-g_{\rho}b_{0}+\sqrt{k_{n}^{2}+(M_{\rm N}^{*})^{2}}\,,$ (3.15)
$\mu_{e}$ $=\sqrt{k_{e}^{2}+m_{e}^{2}}\,,\quad\mu_{\mu}=\sqrt{k_{\mu}^{2}+m_{\mu}^{2}}\,.$ (3.16)
Figure 1: Energy (left panels) and pressure (right panels) densities versus nucleon and DM chemical potentials i.e. $\mu_{\rm N}$ and $\mu_{\chi}$ Each row corresponds to a specific DM particle mass $M_{\chi}=100~{}$MeV (first row), $M_{\chi}=500~{}$MeV (second row), $M_{\chi}=1000~{}$MeV (third row) and $M_{\chi}=5000~{}$MeV (fourth row). We normalize the DM and nucleon chemical potentials with respect to the nucleon mass.
@@
ϵtot=subscriptitalic-$\epsilon_{\text{tot}}=$ $\ \epsilon_{\text{baryons}}+\epsilon_{\text{\rm DM}}+\epsilon_{\text{mesons+Higgs}}+\epsilon_{\text{leptons}}\,,$ (3.18)
ϵbaryons=subscriptitalic-$\epsilon_{\text{baryons}}=$ $\ \frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}d^{3}k\sqrt{k^{2}+(M_{\rm N}^{*})^{2}}\,,$ (3.19)
ϵDM=subscriptitalic-$\epsilon_{\text{\rm DM}}=$ $\ \frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}d^{3}k\sqrt{k^{2}+(M_{\chi}^{*})^{2}}\,,$ (3.20)
ϵmesons+Higgs=subscriptitalic-$\epsilon_{\text{mesons+Higgs}}=$ $\ \frac{1}{2}m_{s}^{2}\phi_{0}^{2}+\frac{1}{3}g_{2}\phi_{0}^{3}+\frac{1}{4}g_{3}\phi_{0}^{4}-\frac{1}{2}m_{V}^{2}V_{0}^{2}-\frac{1}{2}m_{\rho}^{2}b_{0}^{2}+\frac{1}{2}M_{h}^{2}h_{0}^{2}\,,$ (3.21)
[…]

### v2

@@ 1 Introduction
Among others, DM could influence the physical properties and dynamics of compact stars, such as neutron stars, thus leading to a deeper understanding of these objects and the broader dark sector [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Indirect detection possibilities, such as neutrino and gamma-ray signals from inelastic DM annihilations near neutron stars, further increase the potential for uncovering DM properties [39]. Neutron stars can play the role of DM detectors and provide more information about DM properties using the study of the DM interactions with nucleons [26, 40, 36, 41, 42, 43]. Furthermore, the concept of supermassive dark objects broadens insights into DM’s influence on stellar evolution, and white dwarfs mixed with DM may collapse into neutron stars, affecting the overall stellar population [36, 44, 45, 46].
@@
Gravitational wave astronomy is another possible tool to detect DM within neutron stars. Observations from experiments such as LIGO (Laser Interferometer Gravitational-Wave Observatory) and Virgo, including events like GW170817, provide insights into the structure of neutron stars [47, 48]. Non-radial oscillation modes in DM-admixed neutron stars suggest that these oscillations could serve as indicators of DM presence [49, 31, 50, 29, 51]. Additionally, the study of tidal Love numbers–parameters that quantify a star’s tidal response to external gravitational fields–enhances our understanding of the composition of neutron stars [52, 53, 54, 55] and allows for elucidating the possible existence and nature of DM.
@@
Specific DM candidates, including axions, dark photons, and ultralight particles, exhibit unique astrophysical signatures that can be probed through neutron star studies. Axions, for example, may address the cusp-core problem in galactic halos and significantly influence neutron star cooling processes [56, 57, 58]. Dark photons impact neutron star stability and could generate observable signals detectable through various astrophysical observations [59, 60, 42, 44]. Ultralight dark photons and millicharged particles extend DM models with significant implications for neutron star stability and dynamics [59, 60, 61], while studies suggesting neutron decay into a dark Dirac fermion and a dark photon could explain discrepancies in neutron lifetime measurements, offering new pathways for DM detection [42, 43]. Neutron stars can also be used for the detection of DM and probe DM properties through their interactions with nucleons within the core of neutron stars [41, 42, 43]. Exotic compact objects, such as mirror neutron stars and boson stars, offer valuable insights into DM’s astrophysical role. Mirror matter within neutron stars could produce distinctive gravitational wave signatures, differentiating these stars from their ordinary counterparts [62, 63, 4]. Self-interacting DM models, particularly those involving the Higgs portal and muonic forces, are investigated using astrophysical data to verify their viability [64, 65, 66]. Boson stars, formed from bosonic DM with repulsive self-interactions, present alternative compact objects that challenge traditional neutron star models [67, 68, 69, 70, 71, 72]. Studies on fermion-boson stars aim to reconcile observations of compact stars with complex DM interactions, bridging theoretical predictions with astronomical data [10, 68, 73, 74, 75]. Moreover, asymmetric bosonic DM, mediated by vector fields coupled to scalar particles, may alter neutron stars’ mass, radius, and tidal deformability [13, 76, 77, 16, 17, 78, 79]. Non-annihilating particle DM can slowly accumulate within neutron stars, potentially leading to the formation of low-mass black holes, a scenario supported by LIGO-Virgo observations [41, 46, 45, 80, 81, 3]. DM may also influence neutron star heating and cooling, with density-dependent EoS models showing that DM admixtures impact thermal evolution and axion emissions from protoneutron stars suggest possible indirect detection of these particles [77, 57, 58, 82, 83, 84, 85, 86, 87, 88, 89]. These cooling mechanisms, governed by energy loss processes such as neutrino emission, can constrain axion properties and their coupling strengths to ordinary matter, offering a window into physics beyond the Standard Model [82, 57, 58].
@@ 2 Structure Equations: Two Approaches
$\displaystyle\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,$, (2.4)
@@
$\displaystyle\frac{dp(r)}{dr}$ $\displaystyle=$ $\displaystyle-\left(\epsilon(r)+p(r)\right)\frac{d\nu(r)}{dr}\,$, (2.5)
@@
$\displaystyle\frac{dM(r)}{dr}$ $\displaystyle=$ $\displaystyle 4\pi r^{2}\epsilon(r)\,$, (2.6)
@@
$\displaystyle\mu=\frac{\epsilon+p}{\rho}\,$, (2.7)
or, alternatively, it can be derived from the energy density as the derivative with respect to the number density  [^1] For nucleonic matter, one can solve the structure equations based on the baryon chemical potential $\mu_{\rm N}$ by considering the baryon/nucleon number density $\rho_{\rm N}$, which includes both neutrons and protons $\displaystyle\mu_{\rm N}=\frac{\epsilon+p}{\rho_{\rm N}}\,,\quad\mu_{\rm N}=\mu_{n}+\mu_{p}\,,\quad\rho_{\rm N}=\rho_{n}+\rho_{p}\,.$ (2.8) This approach allows for a unified treatment of the matter inside neutron stars, accounting for the contributions from different particle species.
$\displaystyle\mu=\frac{d\epsilon}{d\rho}\,.$ (2.9)
@@
$\displaystyle\nabla_{\mu}T^{\mu\nu}_{\rm N}=-Q^{\nu},\quad\nabla_{\mu}T^{\mu\nu}_{\rm DM}=Q^{\nu}\,$, (2.12)
@@
$\displaystyle\frac{d\mu_{i}(r)}{dr}=-\mu_{i}(r)\frac{d\nu(r)}{dr}\,$, (2.13)
$\displaystyle\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,$, (2.14)
$\displaystyle\frac{dM(r)}{dr}=4\pi r^{2}\epsilon(r)\,$, (2.15)
@@
$\displaystyle\frac{dp_{i}(r)}{dr}=-(\epsilon_{i}(r)+p_{i}(r))\frac{d\nu(r)}{dr}\,$, (2.16)
$\displaystyle\frac{d\nu(r)}{dr}=\frac{G\left(M(r)+4\pi r^{3}p(r)\right)}{r\left(r-2GM(r)\right)}\,.$ (2.17)
$\displaystyle\frac{dM(r)}{dr}=4\pi r^{2}\epsilon(r)\,$, (2.18)
@@ 2.1 Dimensionless structure equations for the interacting DM-nucleonic system
$\displaystyle M^{\prime}=\frac{M}{M_{\rm L}}\,,\quad M_{\rm L}=\frac{M_{\rm Pl}^{3}}{m_{\rm f}^{2}}\,,\quad r^{\prime}=\frac{r}{R_{\rm L}}\,,\quad R_{\rm L}=\frac{M_{\rm Pl}}{m_{\rm f}^{2}}\,,\quad\mu^{\prime}=\frac{\mu}{m_{f}}\,$, (2.19)
@@
$\displaystyle\frac{d\nu(r^{\prime})}{dr^{\prime}}$ $\displaystyle=$ $\displaystyle\frac{\left(M^{\prime}(r^{\prime})+4\pi r^{\prime 3}p^{\prime}(r^{\prime})\right)}{r^{\prime}\left(r^{\prime}-2M^{\prime}(r^{\prime})\right)}\,$, (2.20)
$\displaystyle\frac{dM^{\prime}(r^{\prime})}{dr^{\prime}}$ $\displaystyle=$ $\displaystyle 4\pi r^{\prime 2}\epsilon^{\prime}(r^{\prime})\,$, (2.21)
$\displaystyle\frac{d\mu^{\prime}_{\rm N}(r^{\prime})}{dr^{\prime}}$ $\displaystyle=$ $\displaystyle-\mu^{\prime}_{\rm N}(r^{\prime})\frac{d\nu(r^{\prime})}{dr^{\prime}}\,$, (2.22)
$\displaystyle\frac{d\mu^{\prime}_{\rm DM}(r^{\prime})}{dr^{\prime}}$ $\displaystyle=$ $\displaystyle-\mu^{\prime}_{\rm DM}(r^{\prime})\frac{d\nu(r^{\prime})}{dr^{\prime}}\,.$ (2.23)
@@ 3 The Higgs Portal Fermionic Dark Matter Model
$\displaystyle\mathcal{L}=$ $\displaystyle\ \bar{\psi}\left[\gamma^{\mu}\left(i\partial_{\mu}-g_{v}V_{\mu}-g_{\rho}\tau\cdot b_{\mu}\right)-\left(M_{\rm N}+g_{s}\phi\right)\right]\psi$
$\displaystyle+\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m_{s}^{2}\phi^{2}-\frac{1}{3}g_{2}\phi^{3}-\frac{1}{4}g_{3}\phi^{4}$
$\displaystyle-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}-\frac{1}{4}b_{\mu\nu}b^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}b_{\mu}b^{\mu}\,$, (3.1)
@@
$\displaystyle V_{\mu\nu}$ $\displaystyle=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}\,$, (3.2)
$\displaystyle b_{\mu\nu}$ $\displaystyle=\partial_{\mu}b_{\nu}-\partial_{\nu}b_{\mu}\,.$ (3.3)
@@
$\displaystyle\mathcal{L}_{\text{\rm DM}}=$ $\displaystyle\ \bar{\chi}\left(i\gamma^{\mu}\partial_{\mu}-M_{\chi}+yh\right)\chi+\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}M_{h}^{2}h^{2}+f\frac{M_{\rm N}}{v}\bar{\psi}h\psi\,$, (3.4)
where $h$ is the Higgs field, $y$ is the Yukawa coupling between DM and the Higgs, $f$ represents the nucleon-Higgs form factor  [^2] We fix the parameters to $M_{h}=125$ GeV [113], $y=0.07$ and $f=0.35$ as a reasonable choice, see e.g. Ref. [7]. To assess the sensitivity, varying $y$ or $f$ by a factor of 2 modifies the effective DM mass $M_{\chi}^{*}$ by $\sim 10\%$, leading to small shifts in the mass-radius relation for $M_{\chi}=1000\,\text{MeV}$. Larger deviations could significantly alter the EoS stiffness, a topic we defer to future work. , $M_{\rm N}$ and $M_{\chi}$ are the nucleon and $\chi$ masses, respectively, and $v$ is the vacuum expectation value of the Higgs field. The Higgs portal parameters ($M_{h}=125\,\text{GeV}$, $y=0.07$, $f=0.35$) are chosen based on experimental constraints and theoretical consistency. The Higgs mass is set to its measured value [113], while the Yukawa coupling $y=0.07$ and nucleon-Higgs form factor $f=0.35$ are adopted from prior studies (e.g., [7]) as representative values that ensure weak DM-nucleon coupling without violating direct detection bounds. The couplings we have chosen are small enough to avoid the bounds on Higgs decay from collider data. The overall behavior of the equation of state of the mixed matter has a relatively low sensisivity with respect to the coupling between the two sectors in our set-up. The interaction term $f\frac{M_{\rm N}}{v}\bar{\psi}h\psi$ accounts for the coupling between nucleons and the Higgs field [114]. The RMF model employed here is the NL3. It saturates nuclear matter at $\rho_{0}\approx 0.15\,\text{fm}^{-3}$ with a binding energy of $-16.0\,\text{MeV}$ and supports a maximum neutron star mass of more than 2 $\,M_{\odot}$, aligning with observational constraints. While our RMF model with $\omega$, $\rho$, and $\sigma$ meson interactions (Eq. (3.1)) captures the essential physics of nucleonic matter with Higgs portal DM, richer RMF frameworks exist, such as those incorporating additional meson fields or hyperons [115]. Similarly, alternative dark sector models, e.g., those with different DM interactions [116], could yield distinct effects. Our choice represents a minimal yet thermodynamically consistent extension to explore DM-nucleon coupling via the Higgs portal, serving as a baseline for future comparisons.
@@
$\displaystyle M_{\rm N}^{*}$ $\displaystyle=M_{\rm N}+g_{s}\phi_{0}-\frac{fM_{\rm N}}{v}h_{0}\,.$ (3.6)
As for the mesons, the field equations of motion follow from the Euler-Lagrange equations, where in the mean-field approximation, the meson fields are replaced by their classical expectation values (denoted by the subscript $0$)[7, 108]
$\displaystyle\phi_{0}$ $\displaystyle=\frac{1}{m_{s}^{2}}\left(-g_{s}\langle\bar{\psi}\psi\rangle-g_{2}\phi_{0}^{2}-g_{3}\phi_{0}^{3}\right)\,$, (3.7)
$\displaystyle V_{0}$ $\displaystyle=\frac{g_{v}}{m_{V}^{2}}\langle\psi^{\dagger}\psi\rangle=\frac{g_{v}}{m_{V}^{2}}(\rho_{p}+\rho_{n})\,$,
$\displaystyle b_{0}$ $\displaystyle=\frac{g_{\rho}}{m_{\rho}^{2}}\langle\psi^{\dagger}\tau_{3}\psi\rangle=\frac{g_{\rho}}{m_{\rho}^{2}}(\rho_{p}-\rho_{n})\,$,
$\displaystyle h_{0}$ $\displaystyle=\frac{y\langle\bar{\chi}\chi\rangle+f\frac{M_{\rm N}}{v}\langle\bar{\psi}\psi\rangle}{M_{h}^{2}}\,.$
@@
$\displaystyle\rho$ $\displaystyle=\langle\psi^{\dagger}\psi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}d^{3}k\,$, (3.8)
$\displaystyle\rho_{\rm N}^{s}$ $\displaystyle=\langle\bar{\psi}\psi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}\frac{M_{\rm N}^{*}}{\sqrt{M_{\rm N}^{*2}+k^{2}}}d^{3}k\,$, (3.9)
$\displaystyle\rho_{\text{\rm DM}}^{s}$ $\displaystyle=\langle\bar{\chi}\chi\rangle=\frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}\frac{M_{\chi}^{*}}{\sqrt{M_{\chi}^{*2}+k^{2}}}d^{3}k\,$, (3.10)
@@
$\displaystyle M_{\chi}^{*}$ $\displaystyle=M_{\chi}-yh_{0}\,$, (3.11)
and where $\gamma$ ($=2$) is the spin degeneracy factor for fermions, and $k_{F}$ and $k_{\chi}=k_{F}^{\text{\rm DM}}$ are the Fermi momenta of nucleons and DM particles, respectively  [^3] The number of nucleons can be obtained from the following integral [99]: $\displaystyle N_{\rm N}=\int 4\pi r^{2}\left(1-\frac{2GM(r)}{r}\right)^{-1/2}\rho_{\rm N}(r)\,dr\,.$ (3.12) Similarly, one can estimate the number of DM particles from [99]: $\displaystyle N_{\rm DM}=\int 4\pi r^{2}\left(1-\frac{2GM(r)}{r}\right)^{-1/2}\rho_{\rm DM}(r)\,dr\,.$ (3.13) The fraction of DM particles with respect to nucleons is then given by $f_{\rm DM}={N_{\rm DM}}/{N_{\rm N}}$. . The chemical potentials for nucleons and leptons are given by [99]
$\displaystyle\mu_{p}$ $\displaystyle=g_{v}V_{0}+g_{\rho}b_{0}+\sqrt{k_{p}^{2}+(M_{\rm N}^{*})^{2}}\,$, (3.14)
$\displaystyle\mu_{n}$ $\displaystyle=g_{v}V_{0}-g_{\rho}b_{0}+\sqrt{k_{n}^{2}+(M_{\rm N}^{*})^{2}}\,$, (3.15)
$\displaystyle\mu_{e}$ $\displaystyle=\sqrt{k_{e}^{2}+m_{e}^{2}}\,,\quad\mu_{\mu}=\sqrt{k_{\mu}^{2}+m_{\mu}^{2}}\,.$ (3.16)
Figure 1: Energy (left panels) and pressure (right panels) densities versus nucleon and DM chemical potentials i.e. $\mu_{\rm N}$ and $\mu_{\chi}$ . Each row corresponds to a specific DM particle mass $M_{\chi}=100$~MeV (first row), $M_{\chi}=500$~MeV (second row), $M_{\chi}=1000$~MeV (third row) and $M_{\chi}=5000$~MeV (fourth row). We normalize the DM and nucleon chemical potentials with respect to the nucleon mass.
@@
$\displaystyle\epsilon_{\text{tot}}=$ $\displaystyle\ \epsilon_{\text{baryons}}+\epsilon_{\text{\rm DM}}+\epsilon_{\text{mesons+Higgs}}+\epsilon_{\text{leptons}}\,$, (3.18)
$\displaystyle\epsilon_{\text{baryons}}=$ $\displaystyle\ \frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}}d^{3}k\sqrt{k^{2}+(M_{\rm N}^{*})^{2}}\,$, (3.19)
$\displaystyle\epsilon_{\text{\rm DM}}=$ $\displaystyle\ \frac{\gamma}{(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}d^{3}k\sqrt{k^{2}+(M_{\chi}^{*})^{2}}\,$, (3.20)
$\displaystyle\epsilon_{\text{mesons+Higgs}}=$ $\displaystyle\ \frac{1}{2}m_{s}^{2}\phi_{0}^{2}+\frac{1}{3}g_{2}\phi_{0}^{3}+\frac{1}{4}g_{3}\phi_{0}^{4}-\frac{1}{2}m_{V}^{2}V_{0}^{2}-\frac{1}{2}m_{\rho}^{2}b_{0}^{2}+\frac{1}{2}M_{h}^{2}h_{0}^{2}\,$, (3.21)
$\displaystyle\epsilon_{\text{leptons}}=$ $\displaystyle\ \sum_{l=e,\mu}\frac{1}{\pi^{2}}\int_{0}^{k_{l}}k^{2}\sqrt{k^{2}+m_{l}^{2}}dk\,$, (3.22)